OpenStudy (anonymous):

Find the general solution of: y'+4x=cos2x Can you please show me the steps you took to get the answer of this problem.

myininaya (myininaya):

y'=cos(2x)-4x just integrate both sides to find y

myininaya (myininaya):

\[y=\frac{1}{2}\sin(2x)-4\frac{x^2}{2}+C\]

OpenStudy (anonymous):

dy/dx = cos2x - 4x dy = cos(2x) - 4x dx integrate both sides : y = sin(2x) / 2 - 2x^2 + C

OpenStudy (anonymous):

Oh! you guys are awesome, so every problem like this will involve integrating, right? Okay, I have one more: Find the general solution of: 4xy dx=(x^2+4)dy

myininaya (myininaya):

separate your variables then integrate

myininaya (myininaya):

\[\frac{4x}{x^2+4} dx=\frac{1}{y} dy\]

OpenStudy (jamesj):

I just want to double check on your original problem. It is y'+4x=cos2x NOT y'+4y=cos2x Yes?

OpenStudy (anonymous):

Yes

OpenStudy (jamesj):

great